# Enhanced Computationally Efficient Long LoRA Inspired Perceiver Architectures for Auto-Regressive Language Modeling

### v1

@@ 4.4 Long LoRA Inspired PerceiverAR (LLP) Architecture
The complexity of computing attention is $O(n^{2})$. One way to reduce this complexity is by dividing the input sequence into disjoint segments and only computing the attention in each segment itself. However, the disadvantage of such an approach is that there is information loss due to lack of information flow between segments. LongLoRA Chen et al. solves this problems via shifted sparse attention ($S^{2}$ Attn). In $S^{2}$ Attn, the sequence length is split into different groups and each group computes the attention individually. To support the information flow between different groups, the attention heads are divided in two halves. In the second half of the groups, the tokens are shifted by half the group size. This simple shift causes the information to be shared between neighboring groups. The primary application of LongLoRA was demonstrated in applying LoRA (Low Rank Adaptation) to the self attention layers in extending the sequence or the context length of existing LLMs. For example, LongLoRA can extend the context length of Llama-2 which is originally trained on a sequence length of 4K tokens by finetuning with data using 16K context length. This is accomplished by dividing the context length into 4K sequences. The $S^{2}$-Attn in LongLoRA will divide the attention heads in two groups, such that the first group conducts self-attention from 1st to 4096th tokens. For the second groups, the tokens are shifted by 2048, such that the attention in second group starts from 2049th token and ends at (4096+2028)=6144th token. This approach proposed by LongLoRA does not increase computation complexity, but allows information exchange between different groups. As the information flows through different layers of the Transformer, the information exchange expands to all the segments causing the net attention to become close to the standard full attention.
@@
Inspired by the idea of LongLoRA’s overlapping attention, we apply this concept to enhance the PerceiverAR design by dividing the input sequence into segments, such that the PerceiverAR allows communication of information from previous segment to the current segment. This is accomplished by first dividing the input sequence into disjoint segments. Next each segment is further divided into two halves, a history component and a latent component. The standard PerceiverAR computation is applied to consecutive pairs of half segments. Specifically, $Q$ is computed on the current half segment while computing $K$ and $V$ is done on the current and previous half segments. Thus the PerceiverAR operates upon overlapping segments, with the overlap amount being half the size of a segment. We term this approach as LongLoRA inspired Perceiver (LLP) architecture. It is depicted in Figure 3(d). The ATTN block is a regular attention block which operates upon half of the first segment only. The attention equations governing the ATTN block with segment size $s$ are:
@@
$Output_{atn}=[\text{softmax}(Q_{atn}K^{T}_{atn})V_{atn}]=[A_{atn}V_{atn}]\in\mathbb{R}^{\frac{s}{2}\times d}$ (25)
@@
$Output_{par_{i}}=[\text{softmax}(Q_{par_{i}}K^{T}_{par_{i}})V_{par_{i}}]=[A_{par_{i}}V_{par_{i}}]\in\mathbb{R}^{\frac{s}{2}\times d}$ (29)
@@
where in Equations 26 to 29, $i\in(1,2\cdot s_{total})$ and $s_{total}$ is the total number of segments. Each of the PAR blocks outputs data equal to half of the segment size i.e., $\mathbb{R}^{\frac{s}{2}\times d}$. All blocks use $Q\in\mathbb{R}^{\frac{s}{2}\times d}$ while $K,V\in\mathbb{R}^{s\times d}$ are computed on double the size i.e., on the current half segment and the previous half segment. Thus to implement autoregressive behavior masking is done on the $\frac{s}{2}\times\frac{s}{2}$ part of the attention in PAR block that is of size $\frac{s}{2}\times s$. The very first block in each layer is different and does a normal attention computation with triangular masking on the first half segment. This is done so that even the first half segment can be used in autoregressive implementation, unlike a normal PerceiverAR where the history part cannot be used in autoregressive training in terms of masking this part. All blocks in a layer output data of $\frac{s}{2}\times d$ size. These are then catenated to form an output of size $n\times d$. All layers in the LLP architecture are identical as shown in Figure 3(d).
@@
The main advantage of the proposed LLP architecture is that it provides identical efficiency in terms of computations as the LongLoRA design. Figure 3(c) 3(c) 3(c) shows the sparse attention pattern as it is computed in the first layer, and how each subsequent layer increases the attention receptive field (because of overlap of half segments in the PerceiverAR blocks). As shown in Figure 3(c), after enough layers, the information from all previous segments is available to the PAR block as it calculates the attention on the two half segments. Thus we achieve a similar benefit in calculating the attention as Long LoRA. We also overcome the drawback of the PerceiverAR where the history component of the input sequence cannot be used in autoregressive training. In our LLP model, the entire input sequence can be used in autoregressive training, thus giving us the benefit of highly efficient attention computation, full sequence length auto regressive training, and availability of entire previous context as enough layers are used in the architecture.
@@ 5 Experimental Results
We evaluate different enhanced PerceiverAR architectures on Wikitext-103 Merity et al. (2022) and PG-19 Rae et al. (2019) to test the perplexity of different models. In general, each model is trained under the same set of hyperparameters (batch size, total number of epochs and learning rate scheduler) to make the perplexity comparisons standardized.
Results for Enhanced Architectures V1, V2 and V3 - In Table 1, the main experimental results are presented for three different model configurations (A, B and C). For all configurations, the history length in the Perceiver model is the sequence length subtracted from the latent length. In Configuration A, a sequence length of 1024 is used with an embedding size of 512, with 8 heads and 8 layers. Enhanced Arch V2 uses a segment size of 256 with latent sizes of 256 and 512. It uses a segment size of 128 when the latent length is 768. Enhanced Arch V3 compresses the history component to size 128. In Configuration B, the embedding size is 768, with 6 heads per layer and 6 layers. In this configuration, a sequence length of 2048 is used, with a latent length of 1024. Enhanced Arch V2 uses a segment size of 512. In this configuration, each model is trained for 200,000 iterations with a starting learning rate of $2^{-4}$. Lastly, in Configuration C an embedding size of 768 is used and there are 6 heads per layer with a total of 6 layers. A sequence length of 2048 and a latent length of 1024 is used. Enhanced Arch V2 uses a segment size of 512 and model V3 uses a compression size of 256. Models in this configuration are trained for 500,000 iterations.
From Table 1, it can be seen that Enhanced Arch V1 performs better than other enhanced models, and significantly better than the baseline PerceiverAR. This is due to the fact this enhancement carries the history component to all layers and keeps refining this information in each layer. Unlike the baseline model, where the history information is absorbed into the latent part after the first layer. Enhanced Arch V2 comes close to Enhanced Arch V1’s performance, as it also carries the history information to all layers. However, in Enhanced Arch V2, the history is divided into segments, and the attention is only done within the segment. Computation efficiency wise, this is better than Enhanced Arch V1. Enhanced Arch V3 compresses the history information and uses it in each layer. It does not seem to improve the perplexity (lower is better) significantly as compared to the baseline. Because LLP model is different in its design, we present the results for it separately after comparing the baseline with other model enhancements.
Results for Long LoRA inspired PerciverAR (LLP) model - In Tables 2 and 3, we present results on the Wikitext-103 and PG-19 datasets for the LLP models. We achieve better perplexity than the PercieverAR baseline with significantly smaller (half or less) model size. This can be attributed to the pairwise overlapping application of the PerceiverAR style computation. This results in appropriate information extraction, and propagation of information of the entire context, down the layers of the transformer. Tables 4 and 5 provide a comparison of our LLP model with other SOTA models, including the PercieverAR baseline model, on both the Wikitext-103 and PG-19 datasets. As can be seen from the results, our model achieves the lowest perplexity with the smallest model size.
We test our LLP architecture on the image part of the Long Range Arena (LRA) benchmark Tay et al. . In the image part of LRA, the CIFAR-10 dataset (referred to as sCIFAR-10) is treated as a sequence of grayscale pixels. Table 6 presents the results on image classification for different Transformer architectures. The Transformer architectures compared in Tay et al. used a smaller architecture with 3 layers and 4 heads in each layer. The embedding size used is 64, with a feedforward network size of 128. For fair architecture comparison, we also used the same sizes for one of our LLP models. Our Transformer implementation uses Rotary Position Embedding (RoPE Su et al. (2024)) which yields better accuracy. Even though the state space models perform better than Transformer-based designs on sCIFAR-10, our LLP Transformer model produces the best known accuracy for Transformer-based architectures on the sCIFAR-10 dataset.
@@
Baseline Enhanced Arch V1 Enhanced Arch V2 Enhanced Arch V3
@@
Baseline Enhanced Arch V1 Enhanced Arch V2 Enhanced Arch V3
@@
Baseline Enhanced Arch V1 Enhanced Arch V2 Enhanced Arch V3
@@
LLP Model
Segment Size=512 Segment Size=256 Segment Size=128
Sequence Length=1024 20.2716 20.5148 20.7883
Sequence Length=2048 20.0021 20.4536 20.7482
Table 2: Perplexity results for the LLP model with different segment sizes on Wikitext-103.
Model Architecture LLP - 12 layers LLP - 18 layers LLP - 24 layers PerceiverAR* 60 Layers
Model Size 129.60 million 172.12 million 214.64 million 974.6 million
Test Perplexity Wikitext-103 19.92 17.82 17.43 18.35
Test Perplexity PG-19 21.89 20.42 18.83 28.9
@@
Table 3: Perplexity results on Wikitext-103 and PG-19. All LLP models use 6 heads, an embedding dimension of 768, a sequence length of 2048 and a segment size of 256. PerceiverAR* indicates result cited from Hawthorne et al. (2022).
@@
LLP (ours) 172.12 million 17.82
LLP (ours) 87 million 20.00
@@
Table 4: Comparison of perplexity results on Wikitext-103 with published architectures of similar model sizes. $\dagger$ is from Dai et al. (2019), is from Beck et al. (2024), ** is from Martins et al. (2021) and Δ is from Fu et al. (2022).
@@
LLP Transformer (ours) 24 Layers – 214 million 18.83
Compressive Transformer* 36 Layers – size unknown 33.6
Routing Transformer* 22 Layers – 490 million 33.2
Transformer-XL 36 Layers – size unknown 36.3
Block Recurrent Transformer* 24 Layers – 1.3 billion 26.5
Table 5: Comparison of perplexity results for the LLP model and other models on PG-19 Dataset. * indicates results reported from Hutchins et al. (2022) and $\dagger$ indicates results reported from Dai et al. (2019).
@@
LLP Transformer(ours) – 12 layers 64.42%
LLP Transformer (ours) – 3 layers 59.32%
@@
Table 6: Comparison of Transformer architectures on the sCIFAR-10 dataset. ∗∗∗ indicates results from Tay et al.
6 Discussion of the LLP Algorithm and its Advantages
In the LLP algorithm, the PerceiverAR operation is performed on a pair of overlapping half segments $S_{i}^{j}$, where $i$ is the half segment number and $j$ is the layer number in each layer of Transformer. This is visually shown in Figure 4. The effective receptive attention field increases as computation progresses down the layers of Transformer. For example, the calculation of the output in segment $S4$ in layer 3 is denoted as $S_{4}^{3}$ and uses $S_{4}^{2}$ and $S_{3}^{2}$ from the previous layer (layer 2). $Q$ is computed on $S_{4}^{2}$ while $K$ and $V$ are computed on $S_{3}^{2}$ and $S_{4}^{2}$. Note that $S_{3}^{2}$ in turn uses $S_{2}^{1}$ and $S_{3}^{1}$ from the previous layer (layer 1), and $S_{2}^{1}$ further uses $S_{1}^{0}$ and $S_{2}^{0}$. The segment information accumulated by $S_{4}^{3}$ from previous segments is depicted with a yellow color in Figure 5. Thus, even though the attention computation is local with size $\frac{s}{2}\times s$ if the segment size is $s$, the result is that the propagation of information from all previous segments down the layers occurs. This is due to the fact that PerceiverAR style attention between pair of consecutive overlapping segments is performed in our LLP approach. The first half segment is treated as a special case (as it cannot form a pair with a previous segment), and normal full attention is carried out on it with upper triangular masking, to aid in autoregressive modeling.
Advantages of the LLP Approach - In the LLP approach, efficient attention computation without any loss of contextual information down the layers of the Transformer occurs. This is due to local pair wise segment attention, and overlapping segments in the LLP architecture. One reason, this approach performs better than full attention is that each layer in the early stages is learning to predict the next token with less information. For example, in Figure 4 the predicted tokens in $S_{3}^{1}$ (layer 1) only use the information from $S_{2}^{0}$ and $S_{3}^{0}$ and do not use context from $S_{0}^{0}$ and $S_{1}^{0}$. In layer 2, the computation of $S_{3}^{2}$ has information available from $S_{0}^{1}$, $S_{0}^{2}$ and $S_{0}^{3}$ (but not $S_{0}^{0}$). The following layer (layer 3) will have the information from all previous segments in computation of $S_{3}^{3}$. Thus, there is no loss of information in the next token prediction of any given segment in later layers, but initial layers are learning to predict the next tokens with less information. This is analogous to an implicit dropout in the earlier layers. This may be a contributing factor as to why our LLP model outperforms an equivalent full attention Transformer. In summary, the LLP approach combines efficiency and enhanced learning which results in a significantly improved Transformer architecture, as evidenced by the results on benchmarks.
Figure 4: Visual depiction of the LLP Algorithm indicating increase in the attention receptive field as computation proceeds down the layers of Transformer.

### v2

@@ 5 Efficient Context Propagating Perceiver (ECP) Architecture
One way that attention can be efficiently computed is with LongLoRA [4] which solves this problems via shifted sparse attention ($S^{2}$ Attn). In $S^{2}$ Attn, the sequence length is split into different groups and each group computes the attention individually. To support the information flow between different groups, the attention heads are divided in two halves. In the second half of the groups, the tokens are shifted by half the group size. This simple shift causes the information to be shared between neighboring groups. The primary application of LongLoRA was demonstrated in applying LoRA (Low Rank Adaptation) to the self attention layers in extending the sequence length of existing LLMs.
@@
Inspired by the idea of LongLoRA’s overlapping attention, we apply this concept to enhance the PerceiverAR design by dividing the input sequence into segments, such that the PerceiverAR allows communication of information from previous segment to the current segment. This is accomplished by first dividing the input sequence into disjoint segments. Next each segment is further divided into two halves, a history component and a latent component. The standard PerceiverAR computation is applied to consecutive pairs of half segments. Specifically, $Q$ is computed on the current half segment while $K$ and $V$ are computed on the current and previous half segments. Thus, the PerceiverAR operates upon overlapping segments which allows information flow down the layers. We denote this architecture as the Efficient Context propagating Perceiver (ECP). It is depicted in Figure 5.
Figure 5: Efficient Context propagating Perceiver (ECP)
The ATTN block in Figure 5 is a regular attention block which operates upon half of the first segment only. The attention equations governing the ATTN block with segment size $s$ are:
@@
$Output_{atn}=[\text{softmax}(Q_{atn}K^{T}_{atn})V_{atn}]\\ =[A_{atn}V_{atn}]\in\mathbb{R}^{\frac{s}{2}\times d}$ (25)
@@
$Output_{par_{i}}=[\text{softmax}(Q_{par_{i}}K^{T}_{par_{i}})V_{par_{i}}]\\ =[A_{par_{i}}V_{par_{i}}]\in\mathbb{R}^{\frac{s}{2}\times d}$ (29)
@@
where in Equations 26 to 29, $i\in(1,2\dots s_{total})$ and $s_{total}$ is the total number of segments. Each of the PAR blocks outputs data equal to half of the segment size i.e., $\mathbb{R}^{\frac{s}{2}\times d}$. All blocks use $Q\in\mathbb{R}^{\frac{s}{2}\times d}$ while $K,V\in\mathbb{R}^{s\times d}$ are computed on double the size i.e., on the current half segment and the previous half segment. Thus to implement autoregressive behavior, masking is done on the right $\frac{s}{2}\times\frac{s}{2}$ part of the attention in PAR block. Note that PAR attention is of size $\frac{s}{2}\times s$. The very first block in each layer is different and does a normal attention computation with triangular masking on the first half segment. This is done so that even the first half segment can be used in autoregressive modeling, unlike a normal PerceiverAR where the history part cannot be used in autoregressive training in terms of masking this part. All blocks in a layer output data of $\frac{s}{2}\times d$ size. These are then concatenated to form an output of size $n\times d$. All layers in the ECP architecture are identical as shown in Figure 5.
5.1 ECP Analysis and Advantages
In the ECP algorithm, the PerceiverAR operation is performed on a pair of overlapping half segments $S_{i}^{j}$, where $i$ is the half segment number and $j$ is the layer number in each layer of Transformer. This is visually depicted in Figure 6. The effective receptive attention field increases as computation progresses down the layers of Transformer. For example, the calculation of the output in segment $S4$ in layer 3 is denoted as $S_{4}^{3}$ and uses $S_{4}^{2}$ and $S_{3}^{2}$ from the previous layer (layer 2). $Q$ is computed on $S_{4}^{2}$ while $K$ and $V$ are computed on both $S_{3}^{2}$ and $S_{4}^{2}$. Note that $S_{3}^{2}$ in turn uses $S_{2}^{1}$ and $S_{3}^{1}$ from the previous layer (layer 1), and $S_{2}^{1}$ further uses $S_{1}^{0}$ and $S_{2}^{0}$. The segment information accumulated by $S_{4}^{3}$ from previous segments is depicted with a yellow color in Figure 6. Thus, even though the PerceiverAR style attention computation is local with size $\frac{s}{2}\times s$ if the segment size is $s$, the propagation of information from all previous segments occurs down the layers. The first half segment is treated as a special case (as it cannot form a pair with a previous segment), and normal full attention is carried out on it with upper triangular masking, to aid in autoregressive modeling.
@@ 5.1 ECP Analysis and Advantages
Advantages of the ECP Approach - In the ECP approach, efficient attention computation occurs without any loss of contextual information down the layers of the Transformer. The ECP algorithm provides slightly better efficiency in terms of computations than the LongLoRA design due to PaerceiverAR attention. Figure 1(b-d) shows the sparse attention pattern as it is computed in the first layer, and how each subsequent layer increases the attention receptive field (because of overlap of half segments in the PerceiverAR i.e., PAR blocks). As shown in Figure 1(d), after enough layers, the information from all previous segments is available to the PAR block as it calculates the attention on the two consecutive half segments. Thus we achieve a similar benefit in calculating the attention as LongLoRA with slightly better efficiency, and more flexibility in the number of segments that the input sequence can be partitioned into. Note that the partitioning in LongLoRA is constrained by the number of heads.
One reason our ECP architecture performs better than full attention is that, each layer in the early stages is learning to predict the next token with less information. For example, in Figure 6 the predicted tokens in $S_{3}^{1}$ (layer 1) only use the information from $S_{2}^{0}$ and $S_{3}^{0}$ and do not use context from $S_{0}^{0}$ and $S_{1}^{0}$. In layer 2, the computation of $S_{3}^{2}$ has information available from $S_{0}^{1}$, $S_{0}^{2}$ and $S_{0}^{3}$ (but not $S_{0}^{0}$). The following layer (layer 3) will have the information from all previous segments in computation of $S_{3}^{3}$. Thus, there is no loss of information in the next token prediction of any given segment in later layers, but initial layers are learning to predict the next tokens with less information. This is analogous to an implicit dropout in the earlier layers. This may be a contributing factor as to why our ECP model outperforms an equivalent size full attention Transformer.
Figure 6: Visual depiction of the ECP Algorithm indicating increase in the attention receptive field as computation proceeds down the layers of Transformer.
In summary, the ECP approach combines efficiency and enhanced learning which results in a significantly improved Transformer architecture, as evidenced by the empirical results which we present next.
@@ 6 Experimental Results
Baseline PerceiverAR Double Attention $s$-Split Double Attention Compressed Double Attention
@@
Baseline PerceiverAR Double Attention $s$-Split Double Attention Compressed Double Attention
@@
Baseline PerceiverAR Double Attention $s$-Split Double Attention Compressed Double Attention
@@
Model Architecture ECP - 12 Layers ECP - 18 Layers ECP - 24 Layers PerceiverAR* 60 Layers
Model Size 129.60M 172.12M 214.64M 974.6M
Perplexity Wikitext-103 19.92 17.82 17.43 18.35
Perplexity PG-19 21.89 20.42 18.83 28.9
@@
Table 2: Test Perplexity results on Wikitext-103 and PG-19. All ECP models use 6 heads, an embedding dimension of 768, a sequence length of 2048 and a segment size of 256. PerceiverAR* indicates result cited from [12]. ECP is our model.
We evaluate the three proposed enhanced PerceiverAR architectures and the ECP architecture on Wikitext-103 [18] and PG-19 [19] datasets to test the perplexity of different models. In general, each model is trained under the same set of hyper-parameters (batch size, total number of epochs and learning rate scheduler) to make the perplexity comparisons standardized.
Results for Double Attention, Compressed Double Attention and $s$-Split Double Attention - In Table 1, the main experimental results are presented for three different model configurations. For all configurations, the history length in the Perceiver model is the sequence length subtracted from the latent length. In Configuration A, a sequence length of 1024 is used with an embedding size of 512, with 8 heads and 8 layers. The $s$-Split Double Attention PerceiverAR uses a segment size of $s=256$ with latent sizes of 256 and 512. It uses a segment size of 128 when the latent length is 768. The Compressed Double Attention PerceiverAR compresses the history component to size 128. In Configuration B, the embedding size is 768, with 6 heads per layer and 6 layers. In this configuration, a sequence length of 2048 is used, with a latent length of 1024. The $s$-Split Double Attention PerceiverAR uses a segment size of 512. In Configuration A and B, each model is trained for 200,000 iterations with a starting learning rate of $2\times 10^{-4}$. Lastly, in Configuration C, an embedding size of 768 is used and there are 6 heads per layer with a total of 6 layers. A sequence length of 2048 and a latent length of 1024 is used. The $s$-Split Double Attention PerceiverAR uses a segment size of 512 and the Compressed Double Attention PerceiverAR uses a compression size of 256. Models in this configuration are trained for 500,000 iterations.
From Table 1, it can be seen that the Double Attention PerceiverAR performs better than other PerceiverAR architecture paradigms, and significantly better than the baseline PerceiverAR. In the baseline model, the history information is absorbed into the latent part after the first layer. In contrast, the Double Attention PerceiverAR carries the history component to all layers, and keeps refining this information in each layer. The $s$-Split Double Attention PerceiverAR comes close to Double Attention’s performance, as it also carries the history information to all layers. However, in the $s$-Split Double Attention, the history is divided into disjoint segments, and the attention is only done within each segment. This results in more efficient computation, with some loss of information due to disjoint history segments. The Compressed Double Attention does not seem to improve the perplexity (lower is better) significantly as compared to the baseline. We present the results and analyses for our best performing ECP model next.
Results for Efficient Context Propagating Perceiver - ECP - In Tables 6 and 2, we present results of the ECP models on the Wikitext-103 and PG-19 datasets. ECP achieves better perplexity than the PercieverAR baseline with significantly less computations. This can be attributed to the pairwise overlapping application of the PerceiverAR style computation. This results in appropriate information extraction, and propagation of information of the entire context, down the layers of the transformer. Tables 3 and 4 provide a comparison of our ECP model with other SOTA models, including the PercieverAR baseline model, on both the Wikitext-103 and PG-19 datasets. As can be seen from the results, our model achieves the lowest perplexity with the smallest model size (lower perplexity is better).
We test our ECP architecture on the image part of the Long Range Arena (LRA) benchmark [21]. In the image part of LRA, the CIFAR-10 dataset (referred to as sCIFAR-10) is treated as a sequence of grayscale pixels. Table 5 presents the results on image classification for different Transformer architectures. The Transformer architectures compared in [21] used a smaller architecture with 3 layers and 4 heads in each layer. The embedding size used is 64, with a feedforward network size of 128. For fair architecture comparison, we also used the same sizes for one of our ECP models. Our Transformer implementation uses Rotary Position Embedding (RoPE [20]) which yields better accuracy. Even though the state space models perform better than Transformer-based designs on sCIFAR-10, our ECP Transformer model produces the best known accuracy for Transformer-based architectures on the sCIFAR-10 dataset.
Efficiency and Complexity Analysis - The number of calculation steps needed in the attention for different models is shown in Table 7. Each attention step indicates calculation of one attention entry. For example, in the canonical Transformer, the number of calculation steps in attention will be n × n. As a comparison, the attention steps for different architectures are shown in Table 8 with sequence length n = 4096, layers l = 48, heads h = 24, segment size s = 256 and projection p = 256 . As can be seen from Table 8, the ECP model is extremely efficient with only 12% of the computation needed in attention with respect to full attention in a Transformer. It also performs the best due to its pairwise extraction of attention information and propagation of attention down the layers similar to Long LoRA.
@@
ECP (ours) 172.12 million 17.82
ECP (ours) 87 million 20.00
@@
Table 3: Comparison of perplexity results on Wikitext-103 with published architectures of similar model sizes. $\dagger$ is from [6], is from [2], ** is from [17] and Δ is from [10].
@@
ECP Transformer (ours) 24L (214M) 18.83
Compressive Transformer* 36L (unknown) 33.6
Routing Transformer* 22L (490M) 33.2
Transformer-XL 36L (unknown) 36.3
Block Recurrent Transformer* 24L (1.3B) 26.5
Table 4: Comparison of perplexity results for the ECP model and other models on PG-19 Dataset. * indicates results reported from [13] and $\dagger$ indicates results reported from [6]. L=Number of layers and () gives the number of model parameters in either millions (M) or billions (B).
@@
ECP (ours) – 12 layers 64.42%
ECP (ours) – 3 layers 59.32%
@@
Table 5: Comparison of Transformer architectures on the sCIFAR-10 dataset. ∗∗∗ indicates results from [21]
ECP Architecture
S=512 S=256 S=128
SL=1024 20.2716 20.5148 20.7883
SL=2048 20.0021 20.4536 20.7482
Table 6: Perplexity results for the ECP model with different segment sizes on Wikitext-103. S= Segment Size. SL= Sequence Length.
Model Number of Calculation Steps in Attention
PerceiverAR (baseline) $\left\lceil\frac{n}{2}\times n\right\rceil+(l-1)\left\lceil\frac{n}{2}\right\rceil^{2}\times h$
Double Attention $\left[\left\lceil\frac{n}{2}\times n\right\rceil+\left\lceil\frac{n}{2}\times\frac{n}{2}\right\rceil\right]\times h\times l$
Compressed Attention $\left[\left\lceil\frac{n}{2s}\times\frac{n}{2s}\right\rceil\times s+\left\lceil\frac{n}{2}\times n\right\rceil\right]\times h\times l$
$s$-Split Double Attention $\left[\left\lceil\frac{n}{2}\times n\right\rceil+\left\lceil\frac{n}{2}\times\left(p+\frac{n}{2}\right)\right\rceil\right]\times(l-1)+\left\lceil p\times p\right\rceil\times(l-1)\times h$
ECP Model $\left[\left\lceil\frac{s}{2}\times\frac{s}{2}\right\rceil+\left\lceil\frac{s}{2}\times s\right\rceil\times\left(\left\lceil\frac{n}{s/2}\right\rceil-1\right)\right]\times h\times l$
Regular Transformer $(n\times n)\times h\times l$
Table 7: Attention Complexity in Different Models, n = sequence length, p = projection size, s = segment size, l = layers, h = heads
Model Number of Calculation Steps in Attention Percent of Full Attention
PerceiverAR (baseline) 4932 million 25%
Double Attention 14495 million 75%
Compressed Double Attention 9682 million 50%
$s$-Split Double Attention 14858 million 77%
ECP Attention 2452 million 12%
Transformer (full attention) 19327 million 100%
Table 8: Relative computation efficiency of attention for different models with 24 heads and 48 layers. Sequence Length = 4096.